OpenStudy (anonymous):

1/x+1/y=1/z and xy=z what is the average of x and y?

OpenStudy (anonymous):

average=(x+y)/2. Assuming you wanted it in terms of z, then it's 1/y=1-1/x=(x-1)/x y=x/(x-1) So (x+x/(x-1))/2=avg=(x^2-x+x)/(2x-2)=x^2/(2x-2)

OpenStudy (anonymous):

well, not in terms of z...

OpenStudy (anonymous):

(x+y)/xy=1/z x+y=1 average = 0.5

OpenStudy (anonymous):

1/x+1/y=1/z Solve for y please Thanks